6. In the week before the Superbowl, guacamole cost $5.55/package and 10,255,500 packages were purchased. This week the price has gone down to $4.44/package. How many packages of guacamole will be supplied today if the price elasticity of guacamole supply is 0.5? ANSWER:

7. Assume that advertising shifts the demand curve for Coca-Cola to the right along the supply curve which pushes the Coca-Cola price up by 45%. If the old equilibrium price of Coke was $1.33/liter bottle and the old equilibrium quantity is 13,360.0 million liter bottles, the elasticity of Coca-Cola supply is 0.50 and the elasticity of demand is -1.83, what is the new equilibrium quantity demanded of Coca-Cola? What is the new equilibrium quantity supplied?
Answer #1

6.

the price elasticity of guacamole supply of 0.5 means there will be 50% change in the quantity supplied of the change in the price.

price change=4.44/5.55-1=-20% or 20% decline

change in the quantity supplied expected =0.5*price change=0.5*(-20%)=-10% or 10% decline

How many packages of guacamole will be supplied today=10255500*(1-10%)=9229950 package

the above is answer..

we do only one question based on Chegg rule